OpenStudy (yacoub1993):

Tom has a collection of 30 CDs and Nita has a collection of 18 CDs. Tom is adding 1 CD per month to his collection while Nita is adding 5 CDs per month to her collection. After how many months will Tom and Nita have the same number of CDs? Show your work

OpenStudy (yacoub1993):

i think its 3 months but not sure

OpenStudy (yacoub1993):

@Yttrium

OpenStudy (yttrium):

You're correct bro! It's actually 3. :))

OpenStudy (yacoub1993):

i mean the calculation is correct 30+3 = 18+15 33 = 33

OpenStudy (anonymous):

yaa your write but try to expand ur work

OpenStudy (yacoub1993):

how??

OpenStudy (yttrium):

Let 30+x --> Tom's CD per another additional month 18+5x --> Nita's CD per another additional month 30+x = 18+5x 4x = 12 x = 3. This is the ideal solution for me. @Yacoub1993

OpenStudy (anonymous):

yaa this is a good way but u can also solve it by some othe method

OpenStudy (yacoub1993):

ohhh ok thank you bro